sarah transported 12,500$ in a bank 4 years back. she earned today a total amount of 15,500$ what was the rate of interest charged for her investmeny
1
Expert's answer
2020-09-21T08:57:26-0400

solution


"Initial\ investment, \ P=12,500"

"Final\ value, \ C=15,500"

"Duration, \ n=4\ years"


Let the interest rate "=r"


Assuming her initial investment is compounded annually;



"C= P \ (1+r)^n""15,500 = 12,500 \ (1+r)^4""1.24= (1+r)^4""ln\ (1.24)= 4\ ln\ (1+r)""0.0538= ln\ (1+r)""e^{0.0538}= 1+r""1.05525 -1= 0 .05525=r"

answer: the rate of return on her investment is "5.525\%"